Gary and Larry's age combined adds up to 97. Larry is 5 years older than 3 times Gary's age. How old are they?
Answer by ikleyn(52798)

.
From the condition, you have these two equations


G + L = 97     (1)

L = 3G + 5     (2)


Substitute expression (2) into equation (1) to get


    G + (3G + 5) = 97.


Simplify and solve for G


    4G + 5 = 97

    4G = 97 - 5

    4G = 92

     G = 92/4 = 23.


ANSWER.  Gary is 23 years old.  Larry is  3*23+5 = 74 years old.


CHECK.  23 + 74 = 97.    ! Correct !